OpenStudy (anonymous):

Rico mixed up 500cc of salt solution. He weighed it and found its mass to be 585g. Will his shake-up toy work the way he wants it to if he uses this salt solution? Why or why not? Remember, 1ml=1cc. object of materials and there densities are leaves 1.25g/cc and stars 1.12g/cc. Formula is d=m/v. The is making a toy with plastic stars and leaves. they want to know if you put them in a jar with a salt solution will the stars float and leaves sink. I'm so confused

OpenStudy (anonymous):

objects less dense than the fluid float, objects more dense than a fluid sink. find density of salt solution 585g/500cc = 1.17g/cc leaves density = 1.25g/cc, stars density = 1.12g/cc Stars float, leaves sink. since everything is in g/cc, you dont have to worry about the info 1ml=1cc.

OpenStudy (anonymous):

I still don't know how to put it in the formula (d= m/v).

OpenStudy (anonymous):

density = mass / volume. density of salt solution = 585g(mass)/500cc(volume)